25200.00 needs to be divided in 43.2 shares with 12.5 getting 40% of the money and 31.2 people getting 60%. What is the split?
1
Expert's answer
2018-12-10T10:58:11-0500

If 25200.00 needs to be divided in 43.2 shares with 12.5 getting 40% of the money and 31.2 people getting 60%, then the price per share for 12.5 people is

P1 = 25200*0.4/12.5 = 806.4.

The price per share for other 31.2 people is

P2 = 25200*0.6/31.2 = 484.6.

So, the split ratio is 806.4/484.6 = 1.66 or 1.66:1.